OpenStudy (anonymous):

Consider the equation x^2/9+y^2/9=1, which is the equation of a circle. Explain what you might do in order to change this equation of each of the following: 1. an ellipse 2. a hyperbola 3. a parabola What real life objects are physical models of an ellipse, hyperbola, and parabola?

OpenStudy (anonymous):

an ellipse change one of the denominators so they are not equal

OpenStudy (anonymous):

a hyperbola change the plus sign to a minus sign as well

OpenStudy (anonymous):

\[\frac{x^2}{4}-\frac{y^2}{9}=1\] for example for a hyperbola

OpenStudy (anonymous):

a parabola get rid of one of the squares like \[x^2+y=1\]

OpenStudy (anonymous):

the rest of the question i would do a quick google search planets have elliptical orbits, flashlights have parabolas, but i am just making stuff up